In a 16 km cross country ski race Brett skied the first 7 km at a constant speed and then increased his speed by 3 km/h for the last 9 km. If he finished the entire race in 1 hr. 20 min., what was his speed for the first 7 km?
PLEASE PROVIDE DETAILS ON HOW TO SOLVE AS I NEED TO LEARN!
THANKS,
Lynne

Answer by edjones(7311)

d=distance, s=speed, t=time
d=s*t
let x=speed of 1st 7 km, (x+3)=speed of last 9 km
d/t=s
16/(4/3)=16*(3/4)=12 km/h (average speed for race)
(7x+9(x+3))/16 = average speed for race
(7x+9x+27)/16 =12
(16x+27)/16=12
multiply 16 times both sides: 16x+27=192
subtract 27 from each side: 16x=165
divide 16 into each side: x=10 5/16 km/h Ans.
Ed